Force acting on a particle moving in a straight line varies with the velocity of the particle as F=k/v,where K is a constant.The work done by this force in time t is:

(A)K/v^2.t

(B)2Kt

(C)Kt

(D)2Kt/v^2
1
Expert's answer
2012-12-12T09:52:53-0500
The work is equal to the integral of force multiplied by velocity over time
hence A=int(K/v * v dt) = int(K dt) = Kt

answer is (C)Kt